a nurse is mixing a 3% saline solution and an 8% saline solution to get 2 liters of a 5% saline solution. how many liters of each solution must be combined
Answer by lwsshak3(11628)

a nurse is mixing a 3% saline solution and an 8% saline solution to get 2 liters of a 5% saline solution. how many liters of each solution must be combined
***
let x=amt of 3% solution to mix
2-x=amt of 8% solution to mix
..
3%x+8%(2-x)=5%*2
.03x+.16-.08x=.10
.05x=.06
x=1.2
2-x=0.8
amt of 3% solution to mix=1.2 liters
amt of 8% solution to mix=0.8 liters